OpenStudy (shaik0124):

A train travelling at 72 km/h crosses a platform in 30 sec and a man standing on the platform in 18 sec what is the length of platform? A 240 m B 360 m C 420 m D 600 m

OpenStudy (anonymous):

Use v=x/t

OpenStudy (anonymous):

You'll have to do some unit conversions first

OpenStudy (shaik0124):

(72*5)/18

OpenStudy (anonymous):

As I said, you have to do unit conversions

OpenStudy (anonymous):

Convert 72km/h to m/s

OpenStudy (shaik0124):

x=20m/s

OpenStudy (anonymous):

And no, x represents the distance...the units would be meters.

OpenStudy (shaik0124):

i already made unit conversions 72km/hr=(72*5)/18 m/sec

OpenStudy (anonymous):

Oh I see, yes that's good

OpenStudy (anonymous):

So you have 72 km/h = 20 m/s, what's next?

OpenStudy (shaik0124):

calculate length of platform

OpenStudy (anonymous):

Yeah, so x=vt

OpenStudy (shaik0124):

x=v*t x=20m/s *30 sec

OpenStudy (shaik0124):

x=60 m/s

OpenStudy (anonymous):

20*30 = 600...what's the units of distance?

OpenStudy (anonymous):

The seconds get cancelled out

OpenStudy (shaik0124):

now what to do

OpenStudy (shaik0124):

meters

OpenStudy (shaik0124):

how to calculate length of platform

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\normalsize \text{first u need to find the length of train }\hspace{.33em}\\~\\ \normalsize \text{let the length of train be }D_{train} \hspace{.33em}\\~\\ \normalsize \text{it will be speed times distance needed to cross the man }\hspace{.33em}\\~\\ S_{train}=\dfrac{D_{train}}{T_1} \hspace{.33em}\\~\\ 72~~km/hr=\dfrac{D_{train}}{18~sec} \hspace{.33em}\\~\\ 72\times 10^3~~m/hr=\dfrac{D_{train}}{18\times \dfrac{1}{60\times 60} ~~hr} \hspace{.33em}\\~\\ D_{train}=72\times 10^3 \times \dfrac{18}{60\times 60} ~~metres\hspace{.33em}\\~\\ D_{train}=360 ~~metres\hspace{.33em}\\~\\ \normalsize \text{first u need to find the length of platform }\hspace{.33em}\\~\\ \normalsize \text{let the length of platform be x }\hspace{.33em}\\~\\ \normalsize \text{now to cross the platform the train has to }\hspace{.33em}\\~\\ \normalsize \text{pass the length of train as well as the length of platform, hence }\hspace{.33em}\\~\\ S_{train}=\dfrac{x+D_{train}}{30~~sec} \hspace{.33em}\\~\\ 72\times 10^3~~m/hr=\dfrac{x+360}{\dfrac{30}{60\times 60}~~hrs} \hspace{.33em}\\~\\ \normalsize \text{x will come }x=240~~metres \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):